# Distributed Koopman Operator Learning for Perception and Safe Navigation

### v1

@@ III-B Obstacle prediction and constraint formulation for MPC
The distributed Koopman learning algorithm introduced earlier serves as the foundation for prediction-based obstacle avoidance.
To clarify the Koopman prediction process, the paired matrices $X$ and $Y$ encode the temporal transitions of the spatial density field observed by the robot or aerial sensor. Once the distributed Koopman operator $K$ is estimated, it serves as a linear propagator:
@@ IV Simulation Case-Study
The following four-stage simulation study demonstrates and investigates the full perception-to-control pipeline of our method.
@@ IV-A Evaluation of distributed Koopman operator learning
The evaluation begins by characterizing the sensory snapshots used as input to the distributed Koopman learning algorithm. Each snapshot represents a two-dimensional grid of the environment, where entries take values in $[0,1]$ and encode normalized occupancy intensities: regions with values near $1$ correspond to dense obstacle areas, while values close to 0 indicate free space. Thus, each snapshot compactly captures the spatial distribution of dynamic obstacles at a specific time, forming a high-dimensional state of the environment. Sequential snapshots collected over time constitute a temporal dataset describing the spatiotemporal evolution of moving obstacles. In the simulations, dynamic obstacle fields are generated over a $30\times 30$ spatial grid, observed through top-down camera snapshots recorded across $N=10$ time steps. Each frame depicts multiple moving obstacles with distinct initial densities and heterogeneous motion patterns, forming the dataset for distributed Koopman operator learning and subsequent prediction analysis.
Figure 1: Prediction discrepancy $O(t)$ between distributed and centralized Koopman operators over iterations.
@@
The task is to reconstruct and predict the spatiotemporal evolution using $p=3$ nodes connected via a circular undirected communication graph. Each node holds local data matrices $X_{i},Y_{i}$, as portions of the centralized matrices $X,Y$ (with $n=900$ from grid vectorization). Then, each node estimates a local Koopman operator $K_{i}\in\mathbb{R}^{n\times n_{i}},~n_{i}=n/p$, collectively approximating $K\in\mathbb{R}^{n\times n}$, via $100$ optimization steps with neighbor communication. Using matrix $\mathbf{M}$ from (17), we compute $\alpha_{\max}=0.2095$ from (28) and set $\alpha=0.5\alpha_{\max}$, yielding $\rho_{\max}=0.9185$ by (30).
$O(t):=\left\|(K^{*}-K_{\mathrm{d}}{(t)})X\right\|_{F}.$
Here, $K^{*}$ denotes the centralized Koopman operator derived from (13) and $K_{\mathrm{d}}(t)$ the distributed estimate at iteration $t$. The metric quantifies the difference in predicted lifted dynamics. As shown, $O(t)$ decays exponentially, confirming convergence to the centralized solution. The convergence rate depends on the step size, graph Laplacian spectrum, and matrix $\bm{X}$, with $\rho_{\max}=0.9185$, validating the effectiveness of the distributed learning algorithm. Fig. 2 shows the entrywise absolute difference $K_{\Delta}:=|K_{\mathrm{d}}-K^{*}|$, with $K_{\Delta}]_{ij}:=|[K_{\mathrm{d}}]_{ij}-[K^{*}]_{ij}|.$ The heatmap reveals that most entries have small errors, indicating strong alignment between distributed and centralized models. Block patterns reflect the partitioning across $p=3$ nodes, and higher deviations appear sparsely, likely due to local variations or communication limits. Fig. 3 compares eigenvalues of $K^{*}$ (blue) and $K_{\mathrm{d}}$ (red) in the complex plane. All lie within the unit circle, confirming stability. $K_{\mathrm{d}}$’s spectrum is slightly more spread but closely matches dominant modes of $K^{*}$, indicating successful preservation of key system dynamics via distributed learning. Fig. 4 depicts the prediction error of $K_{\mathrm{d}}$ over a $30\times 30$ spatial grid and prediction steps. Errors grow over time due to recursive forecasting, with some localized spikes likely from decentralized update limitations. This spatial-temporal map complements global metrics by revealing when and where prediction errors concentrate.
In the simulations, the lifting function is realized through direct vectorization of the spatial grid, where each pixel intensity of the snapshot is treated as an observable. This straightforward and non-parametric embedding ensures that the lifted state space coincides with the perceptual domain, preserving all spatial information prior to Koopman learning. Such direct lifting is particularly suitable for dynamic environments with unknown structure, as it avoids biases introduced by handcrafted features or predefined basis functions.
Remark 1.
While alternative liftings (such as polynomial bases, radial or Fourier expansions, and neural observables) can improve representation capacity, they often increase design complexity and computational cost. Direct vectorization, by contrast, provides a fully data-driven baseline that captures the complete spatiotemporal variability in the observations, enabling a clear assessment of the scalability and convergence of the proposed distributed Koopman learning algorithm. Furthermore, since density values are already represented on the same grid, this lifting aligns naturally with the Gaussian mixture and polytope-based obstacle extraction pipeline, ensuring seamless integration between perception, learning, and prediction. For occupancy-like fields, direct vectorization yields a locality-preserving and interpretable lifting: each observable corresponds to a fixed spatial location, allowing the Koopman operator to encode transport and deformation patterns across neighboring pixels. This approach eliminates basis-mismatch issues and hyperparameter tuning required by learned dictionaries or kernels, providing a lightweight, transparent, and computationally predictable embedding which results in well suited online updates and integration within the MPC loop.
Remark 2.
@@ IV-B Comparison with other distributed Koopman methods
This subsection benchmarks the proposed distributed Koopman learning algorithm against representative methods from the literature to evaluate its performance under identical experimental conditions. For a consistent comparison, the same dataset, lifting function, and data partitions described earlier were used for the implementations in [liu2020towards] and [nandanoori2021data]. All methods operated on identical splits of the lifted matrices $X$ and $Y$, and were evaluated using normalized one-step and multi-step trajectory prediction errors.
Results indicate that all distributed estimators closely approximate the centralized Koopman operator, confirming the general consistency of distributed EDMD formulations. To quantify performance, we define the normalized one-step and multi-step prediction errors as
@@
where $Y^{(h)}$ represents the $h$-step-ahead lifted data. These normalized quantities measure how closely the distributed approximation replicates the centralized Koopman model at different prediction horizons.
Numerical evaluations show that the one-step prediction errors remain close to unity for all distributed methods ($e_{1}\approx 1.01{-}1.02$), indicating that all the proposed approaches reproduce the centralized Koopman predictions with only a marginal 1–2% deviation. For longer prediction horizons ($h=14$, corresponding to the MPC planning window in the following section), we evaluate the relative performance among the distributed approaches. In this case, our proposed method achieves a normalized multi-step error approximately $11\%-13\%$ smaller than those in [liu2020towards] and [nandanoori2021data]. This improvement reflects enhanced stability and reduced error accumulation when repeatedly applying the learned operator for multi-step forecasts. From a control standpoint, this superior multi-step accuracy is especially valuable: accurate long-horizon Koopman predictions enable the MPC to anticipate dynamic obstacle evolution more reliably, reducing constraint violations and ensuring smoother, safer navigation. Hence, even modest improvements in prediction consistency translate into tangible performance gains for the overall perception–control pipeline.
@@ IV-C MPC with Koopman-based dynamic obstacle avoidance
After validating the accuracy and convergence of the distributed Koopman operator in modeling dynamic obstacles, we integrate it into an MPC framework for safe robot navigation. The Koopman-based forecasts are exploited to construct time-varying polytopic constraints that guarantee collision-free trajectories over the prediction horizon while driving the robot toward its goal through smooth, feasible control actions.
@@
Fig. 5 shows the closed-loop trajectory in an environment with $L_{o}=12$ moving obstacle clusters where we observe that the robot reaches its target smoothly, maintaining safe separation by continuously adapting to predicted obstacle motion.
To further highlight predictive avoidance, Fig. 6 displays key snapshots during navigation. Each depicts the robot’s orientation, predicted trajectory, and evolving obstacle regions, confirming proactive maneuvering guided by Koopman-based forecasts. The robot adjusts its path before collisions become imminent, demonstrating anticipatory control. Fig. 7 quantifies safety margins by plotting the Euclidean distance between the robot and each obstacle center. All distance profiles remain above the black dashed safety threshold, verifying consistent clearance and robust avoidance. This reinforces the predictive and conservative nature of the MPC policy. Lastly, Fig. 8 presents the control input profiles. The accelerations $(a_{x},a_{y})$ govern translational dynamics, while the angular velocity $\omega$ adjusts heading. These signals confirm dynamically feasible, well-coordinated actuation, validating that safe navigation is achieved through efficient integration of Koopman-based predictions within the MPC framework. For a detailed implementation guideline of Koopman-based moving obstacle avoidance MPC, the reader is referred to our recent work [bueno2025koopman], where a complete prediction and control pipeline is developed and validated through a ROS2–Gazebo implementation.
@@
The scalability of the proposed framework stems from both its distributed Koopman learning and its MPC integration. During learning, the lifted data matrices are row-partitioned so that each agent $i\in\mathcal{P}$ processes only local submatrices $X_{i}$ and $Y_{i}$. This partitioning reduces per-agent memory to $\mathcal{O}(n_{i}N)$, compared to $\mathcal{O}(nN)$ centrally, and bounds the computational cost per iteration to $\mathcal{O}(n_{i}^{2}N)$, which remains constant under weak scaling as the number of agents $p$ increases. Communication overhead is minimal since the update rules in (14) exchange only compact summary variables ($S_{i}$ and implicitly $K_{i}$) among neighbors. Consequently, bandwidth requirements depend on the graph connectivity rather than data dimension or sample size. Convergence is governed by the contraction factor $\rho_{\max}$, ensuring that the required iteration count grows only mildly with network size, enabling near-linear wall-clock acceleration with additional agents. In the MPC stage, computational efficiency is maintained when embedding Koopman-based obstacle forecasts into linear constraints. Each predicted obstacle is represented by $n_{\varsigma}$} supporting half-spaces, with ${n_{\varsigma}}=8$ providing a balanced trade-off between precision and complexity. Enforcing only the most active hyperplane per obstacle and prediction step reduces the number of constraints from $\mathcal{O}(L_{o}H{n_{\varsigma}})$ to $\mathcal{O}(L_{o}H)$ while maintaining conservative safety margins. Because these constraints are linear and preserve the sparsity structure of the nominal dynamics (4), the resulting MPC problems remain tractable and can be efficiently warm-started across iterations. Overall, the framework achieves scalable, real-time operation suitable for high-dimensional, dynamic environments.
@@
IV-D Linking Koopman predictions to MPC
constraints via intensity–polytope architecture
To make explicit how perception feeds planning, we visualize the predicted spatial intensities $\rho_{t+h}$ produced by the distributed Koopman operator and the geometric obstacle sets used by MPC at representative time that coincide with Fig. 6(a). At time $t$, the most recent lifted snapshot is propagated forward as $X_{t+h}\approx K^{h}X_{t}$, $h=1,\dots,H$, then mapped back to the grid to obtain $\rho_{t+h}\in[0,1]$. For each predicted map, a threshold $c_{\rho}$ selects high-likelihood occupancy $\Psi_{t}=\left\{\xi\in\mathbb{R}^{2}:\rho_{t}(\xi)>c_{\rho}\right\}$, to which Gaussian mixture model is fit, yielding centers $\{\mu_{\ell,t+h}\}$ and covariances $\{\Sigma_{\ell,t+h}\}$. Confidence ellipses derived from $(\mu_{\ell,t+h},\Sigma_{\ell,t+h})$ are then approximated by polygons and inflated by the safety margin $\varepsilon$, producing time-varying polytopes that circumscribe the predicted obstacle mass. Fig. 9 overlays the predicted intensity field $\rho_{t+h},~t=15,~h=1,7,14$ to show intensity values inside the union of all nominal polytopes (white background elsewhere) that are passed to MPC as the linear constraints in (III-B). This view reveals three key connections: (a) the distributed Koopman operator forecasts the future spatial distribution of moving obstacles on the original grid, (b) the Gaussian mixture model serves as a compact surrogate that locates and shapes obstacle components directly from $\rho_{t}$, and (c) the polygonal circumscribers translate these probabilistic predictions into tractable, conservative linear inequalities enforceable at each step of the MPC horizon. Panels at $t=15$ shown in Fig. 9 align with the controller’s receding-horizon updates used in Fig. 6(a) for $h=1,7,14$. As time progresses, the Koopman predictions advect and deform the intensity ridges, while the Gaussian mixture model components smoothly re-center and re-orient. The resulting polytopes track these changes with ${n_{\varsigma}}=8$ facets, and the inflated boundaries define the actual constraints enforced in (6d). This predictive pipeline explains the anticipatory deviations of the robot path: the controller looks up to the future occupancy through $\rho_{t+h}$, receives a compact linear description via the polytopes, and adjusts the trajectory before high-density regions intersect the planned path.

### v2

@@ III-B Obstacle prediction and constraint formulation for MPC
The distributed Koopman learning algorithm introduced earlier serves as the foundation for prediction-based obstacle avoidance. To clarify the Koopman prediction process, the paired matrices $X$ and $Y$ encode the temporal transitions of the spatial density field observed by the robot or aerial sensor. Once the distributed Koopman operator $K$ is estimated, it serves as a linear propagator:
@@ IV Simulation Case-Study
A graphical overview of the simulated perception-learning-control scenario is provided in Fig. 1. The setup consists of $p=3$ UAV agents collaboratively monitoring a planar area of dimension $q=q_{x}q_{y}$, corresponding to the state space of the unknown nonlinear dynamics in (6). The UAVs are connected through a circular undirected communication graph and are color-coded for clarity. Each UAV observes only a local field of view, indicated by the highlighted ground grids with matching colors. In particular, the field of view of agent $i$ has dimension $q_{i}=q_{x,i}q_{y,i}$; in the simulations we set $q_{x}=q_{y}=30$, yielding $q=900$, and choose an equal spatial partition with $q_{1}=q_{2}=q_{3}=300$.
The monitored area contains multiple moving robot-like obstacles, depicted in black, each following the direction indicated by dashed arrows. The sensory snapshots collected by the UAVs are represented as two-dimensional occupancy grids with values in $[0,1]$, where values close to $1$ indicate dense obstacle regions and values near 0 correspond to free space. Sequential snapshots collected over time form a spatiotemporal dataset capturing the evolution of these dynamic obstacles. In the simulations, each UAV records local snapshots over $N=10$ time steps using onboard camera measurements; the corresponding intensity color bar is shown in Fig. 1.
To enable distributed Koopman learning, local grid vectorization is employed as the lifting function, so that $n_{i}=q_{i}$ for all agents. In parallel, a nonholonomic ground robot (shown in green) navigates the environment toward a target location indicated by a green circle at the center of the grid. The dotted green arrows illustrate candidate trajectories generated by the MPC scheme. Predictions obtained from the distributed learning of obstacle motion are used to forecast future obstacle density maps, which are approximated by Gaussian mixture models and converted into convex polytopic regions. These polytopes enter the MPC formulation as time-varying linear constraints, enabling anticipatory and collision-free motion planning in a dynamic environment.
Figure 1: Graphical illustration of the distributed perception-learning-control scenario.
Figure 2: Prediction discrepancy $\mathcal{O}(t)$ between distributed and centralized Koopman operators over iterations.
@@ IV-A Evaluation of distributed Koopman operator learning
The distributed Koopman learning algorithm is executed for $200$ optimization steps with neighbor communication. Using the matrix $\mathbf{M}$ defined in (17), we compute $\alpha_{\max}=0.2095$ and select the stepsize as $\alpha=0.5\,\alpha_{\max}$, yielding a contraction factor $\rho_{\max}=0.9185$ according to (30). As discussed in Remark 1, this step-size can be chosen in a distributed manner based on global bounds obtained via local computations and consensus, and can be shared consistently by all agents throughout the learning process.
$\mathcal{O}(t):=\left\|(K^{*}-K_{\mathrm{d}}{(t)})X\right\|_{F}.$ (33)
Here, $K^{*}$ denotes the centralized Koopman operator derived from (12) and $K_{\mathrm{d}}(t)$ the distributed estimate at iteration $t$. The metric quantifies the difference in predicted lifted dynamics. As shown, $\mathcal{O}(t)$ decays exponentially, confirming convergence to the centralized solution. The convergence rate depends on the step size, graph Laplacian spectrum, and matrix $\bm{X}$, with $\rho_{\max}$, validating the effectiveness of the distributed learning algorithm. Fig. 3 shows the entrywise absolute difference $K_{\Delta}:=|K_{\mathrm{d}}-K^{*}|$, with $K_{\Delta}]_{ij}:=|[K_{\mathrm{d}}]_{ij}-[K^{*}]_{ij}|.$ The heatmap reveals that most entries have small errors, indicating strong alignment between distributed and centralized models. Block patterns reflect the partitioning across $p=3$ agents, and higher deviations appear sparsely, likely due to local variations or communication limits. Fig. 4 compares eigenvalues of $K^{*}$ (blue) and $K_{\mathrm{d}}$ (red) in the complex plane. All lie within the unit circle, confirming stability. $K_{\mathrm{d}}$’s spectrum is slightly more spread but closely matches dominant modes of $K^{*}$, indicating successful preservation of key system dynamics via distributed learning. Fig. 5 depicts the prediction error of $K_{\mathrm{d}}$ over the whole spatial grid and prediction steps. Errors grow over time due to recursive forecasting, with some localized spikes likely from decentralized update limitations. This spatial-temporal map complements global metrics by revealing when and where prediction errors concentrate.
@@
In the simulations, the lifting is implemented via direct vectorization of the spatial grid, treating each pixel intensity as an observable. This non-parametric choice preserves all spatial information by aligning the lifted space with the perceptual domain, making it well suited for dynamic environments with unknown structure. To assess the impact of lifting richness, we consider the following two performance metrics:
$\mathcal{Q}_{1}:=\frac{\|Y-K_{\mathrm{d}}(t_{\max})X\|_{F}}{\|Y\|_{F}},~\mathcal{Q}_{2}:=\frac{\|K^{*}-K_{\mathrm{d}}(t_{\max})\|_{F}}{\|K^{*}\|_{F}},$
in addition to the transient prediction discrepancy $\mathcal{O}(t)$ defined in (33). For direct vectorization of local raw data segments, $\mathcal{O}(t)$ reaches $10^{-3}$ at approximately $163$ iterations, with $\mathcal{Q}_{1}$ on the order of $10^{-5}$ and $\mathcal{Q}_{2}\simeq 0.43$. Augmenting the local lifting with first- and second-order powers yields comparable accuracy, with $\mathcal{O}(t)$ reaching $10^{-3}$ at around $170$ iterations, $\mathcal{Q}_{1}$ remaining on the order of $10^{-5}$, and $\mathcal{Q}_{2}\simeq 0.50$. Including up to third-order powers further increases the lifted dimension but does not improve performance: $\mathcal{O}(t)$ reaches $10^{-3}$ after about $174$ iterations, while $\mathcal{Q}_{1}$ remains at the same order and $\mathcal{Q}_{2}$ increases to approximately $0.55$. These results indicate that, in this setting, enriching the lifting does not necessarily improve either prediction accuracy or convergence speed and may even degrade agreement with the centralized solution, while incurring higher per-iteration computational cost. This highlights that richer liftings are not universally beneficial and that lightweight, well-aligned representations can offer a more effective accuracy–efficiency trade-off in distributed Koopman learning.
@@
Remark 4.
In practical sensing scenarios, the fields of view of different UAV agents may partially overlap, leading to repeated measurements of the same physical grid cells across agents. Such overlap does not invalidate the distributed Koopman learning algorithm: convergence of Algorithm 1 guarantees recovery of the optimal Koopman operator associated with the chosen lifted coordinate system, even if this representation contains redundant (duplicated) states. However, convergence alone does not automatically remove redundancy; if overlapping grid cells are included multiple times, the learned operator corresponds to a redundant but dynamically consistent coordinate representation. In practice, redundancy can be handled in several standard and fully valid ways. A common approach is to impose a unique global indexing of spatial cells prior to learning, assigning each physical state to a single agent and thereby avoiding duplication altogether. Alternatively, overlapping coordinates may be retained during learning and subsequently fused through a linear aggregation or averaging map, yielding a reduced Koopman model defined on a unique global state space. Such post-processing steps preserve the learned dynamics while eliminating redundancy when a minimal representation is desired. In the presented simulations, a non-overlapping spatial partition is adopted for clarity. Nevertheless, the proposed distributed framework naturally accommodates overlapping fields of view through the above mechanisms, making it directly applicable to realistic multi-UAV sensing configurations.
@@ IV-B Comparison with other distributed Koopman methods
The proposed distributed Koopman learning algorithm is benchmarked against representative approaches from the literature under identical experimental conditions. For a fair comparison, the same dataset, lifting function, and data partitions were used for the methods in [liu2020towards] and [nandanoori2021data]. All algorithms operated on identical splits of the lifted data matrices $X$ and $Y$ and were evaluated using normalized one-step and multi-step prediction errors.
The normalized prediction metrics are defined as
@@
where $Y^{(h)}$ denotes the $h$-step-ahead lifted data. Numerical results show that all distributed methods closely approximate the centralized Koopman operator, with comparable one-step errors ($e_{1}\approx 1.01{-}1.02$). For longer horizons ($h=14$), corresponding to the MPC planning window, the proposed method achieves a $11\%-13\%$ reduction in multi-step prediction error relative to [liu2020towards, nandanoori2021data]. This improved long-horizon accuracy enhances the reliability of Koopman-based predictions for MPC, yielding smoother and safer navigation in dynamic environments.
@@ IV-C MPC with Koopman-based dynamic obstacle avoidance
After validating the distributed Koopman model, its predictions are embedded into an MPC framework as time-varying polytopic constraints, enabling safe and goal-directed robot navigation over the prediction horizon.
@@
Fig. 6 shows the closed-loop trajectory in an environment with $L_{o}=12$ moving obstacle clusters where we observe that the robot reaches its target smoothly, maintaining safe separation by continuously adapting to predicted obstacle motion. Fig. 7 illustrates representative navigation snapshots, showing the robot’s orientation, predicted trajectory, and evolving obstacle regions, and confirming anticipatory maneuvering enabled by Koopman-based forecasts. Fig. 8 reports the Euclidean distance between the robot and obstacle centers, which consistently remains above the safety threshold, verifying robust collision avoidance. Fig. 9 shows the corresponding control inputs, where translational accelerations $(a_{x},a_{y})$ and angular velocity $\omega$ yield dynamically feasible and coordinated motion. Together, these results demonstrate safe and predictive navigation achieved through integration of Koopman-based predictions into MPC. For implementation details, see [bueno2025koopman].
@@
Scalability of the proposed framework arises from both distributed Koopman learning and its MPC integration. During learning, row-wise partitioning of the lifted data ensures that each agent processes only local blocks $X_{i},Y_{i}$, reducing per-agent memory to ${O}(n_{i}N)$ and computation to ${O}(n_{i}^{2}N)$ per iteration, which remains constant under weak scaling as the number of agents grows. Communication overhead is limited to exchanging compact variables ($S_{i}$), making bandwidth requirements dependent on graph connectivity rather than data dimension. Convergence is governed by the contraction factor $\rho_{\max}$, leading to near-linear wall-clock speedup with additional agents. In the MPC stage, Koopman-based obstacle predictions are embedded as linear constraints using polygonal approximations with ${n_{\varsigma}}=8$ facets. Activating only the most restrictive hyperplane per obstacle and prediction step reduces constraint complexity from ${O}(L_{o}Hn_{\varsigma})$ to ${O}(L_{o}H)$ while preserving conservative safety margins. The resulting sparse linear constraints maintain MPC tractability and enable efficient warm-starting, supporting real-time operation in high-dimensional dynamic environments.
@@
IV-D Linking Koopman predictions to MPC constraints
To clarify how perception informs planning, we visualize the predicted spatial intensities $\rho_{t+h}$ generated by the distributed Koopman operator together with the geometric obstacle sets used by MPC at representative times corresponding to Fig. 7(a). From the most recent lifted snapshot at time $t$, predictions are obtained via $X_{t+h}\approx K^{h}X_{t}$, $h=1,\dots,H$, and mapped back to the grid to recover $\rho_{t+h}\in[0,1]$. A threshold $c_{\rho}$ selects high-likelihood regions $\Psi_{t}=\{\xi\in\mathbb{R}^{2}:\rho_{t}(\xi)>c_{\rho}\}$, to which a Gaussian mixture model is fitted, yielding component means $\{\mu_{\ell,t+h}\}$ and covariances $\{\Sigma_{\ell,t+h}\}$. Confidence ellipses derived from $(\mu_{\ell,t+h},\Sigma_{\ell,t+h})$ are then polygonally approximated and inflated by a safety margin $\varepsilon$, producing time-varying polytopes that conservatively bound the predicted obstacle mass. Fig. 10 overlays $\rho_{t+h}$ at $t=15$ for $h=1,7,14$, highlighting intensity values within the union of all nominal polytopes, which are passed to MPC as the linear constraints in (III-B). This visualization illustrates three key links: (i) the Koopman operator predicts future obstacle distributions on the original grid, (ii) the Gaussian mixture model provides a compact parametric surrogate of $\rho_{t}$, and (iii) polygonal circumscribers convert these probabilistic predictions into conservative linear inequalities enforceable along the MPC horizon. As time evolves, the intensity ridges advect and deform, the mixture components re-center and re-orient, and the resulting polytopes, constructed with $n_{\varsigma}=8$ facets and inflated boundaries, define the actual constraints enforced in (13d). This pipeline explains the anticipatory behavior of the robot: future occupancy is forecast through $\rho_{t+h}$, summarized by polytopes, and incorporated into MPC before high-density regions intersect the planned path.